Circular No. 6409

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996ae IN NGC 5775 A. Vagnozzi, G. Piermarini, and V. Russo, S. Lucia Observatory, report

their discovery of an apparent supernova (mag 16.5) in NGC 5775, found on a CCD image exposed on May 21.951 UT; they provide the following position for the new star: α = 14^{h}53^{m}59^{s}.81, δ = +3^{o}31′46′′.3 (equinox 2000.0), which is ∼ 31′′ east and 54′′ south of the center of NGC 5775. Confirmatory images (limiting stellar mag ∼ 19.5) were taken at S. Lucia Observatory on May 21.969, 21.976, 22.903, and 23.939, and by A. Nakamura at Kuma, Ehime, Japan, on May 24.649. G. Hurst, The Astronomer, reports that the object is also present on CCD images obtained on May 25.9–26.0 near mag ∼ 17 by M. Armstrong (Rolvenden, U.K.), D. Strange (Worth Maltravers, U.K.), and T. Platt (Binfield, U.K.); the new star is not present on the Carnegie Atlas of Galaxies, panel 289.

MXB 1730–335 W. H. G. Lewin, E. Morgan, D. W. Fox, and R. Rutledge, Mas-

sachusetts Institute of Technology; L. Bildsten, University of California, Berkeley; T. Dotani, Institute of Space and Astronautical Science; L. Lubin, Carnegie Observatories; M. van der Klis, Astronomical Institute ‘Anton Pannekoek’, University of Amsterdam (UA); and J. van Paradijs, University of Alabama in Huntsville and UA, report: “Following detection of the rapid burster MXB 1730–335 with the RXTE satellite’s All-Sky Monitor (IAUC 6390), we observed the source on May 3, 7, and 13 with its PCA (2– 60 keV). On the first two occasions, it exhibited type-II bursts (spasmodic accretion) with durations of 10–15 s and recurrence times of roughly 100 s (May 3) and 400 s (May 7). These bursts exhibited the usual ‘ringing’ during their decay (cf. Tan et al. 1991, MNRAS 251, 1). Detailed inspection of the burst power spectra also shows 5- to 7-Hz quasi-periodic oscillations on May 3 but not on May 7. No coherent power is seen above the 99.9-percent confidence threshold between 10 and 2048 Hz on either date. For the first two observations, we can place an upper limit of 44 mCrab on any persistent emission from the rapid burster (assuming a Crab-like spectrum). On the third observation, no bursts were observed, and our upper limit for any persistent emission is 21 mCrab. Quick-look results from the ASM team indicate that the source is no longer active.”

1996 May 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green